From a committee consisting of 5 men and 6 women, a sub-committee is formed consisting of 4 men and 3 women. How many different subcommittees are possible?
Answer by greenestamps(13208)



The number of ways of choosing 4 of the 5 men is C(5,4) "5 choose 4" = 5.

The number of ways of choosing 3 of the 6 women is C(6,3) "6 choose 3" = (6*5*4)/(3*2*1) = 20.

The number of ways of choosing 4 of th5 men AND 3 of the 6 women is C(5,4) times C(6,3) = 5*20 = 100.